OpenStudy (anonymous):

The price of 9-volt batteries is increasing according to the function below, where t is years after January 1, 1980. During what year will the price reach $4? P(t) = 1.1 e^(0.047t)

OpenStudy (anonymous):

4=1.1 e^(0.047t)

OpenStudy (anonymous):

divide both side by 1.1

OpenStudy (anonymous):

4/1.1=e^(0.047t)

OpenStudy (anonymous):

ln (4/1.1)=0.047t

OpenStudy (anonymous):

so wats t?

OpenStudy (anonymous):

for you to solve

OpenStudy (anonymous):

so do i divide ln(4/1.1) by 0.047?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

27.46 is that correct?

OpenStudy (anonymous):

but the question is asking what year will the price reach $4